# On the composable security of weak coin flipping

### v2

@@ 2.4 Weak coin flipping
Here we formulate WCF protocols with causal boxes.
@@
A weak coin flipping protocol using noiseless communication channels is a pair of converters $\pi=(\pi_{A},\pi_{B})$ with compatible inner interfaces. $\pi_{A}$ ($\pi_{B}$) has one output wire on the outer interface and only outputs a single classical bit at $t_{a}$ ($t_{b}$). The poset $\mathcal{T}$ of the protocol is bounded both from above and below, i.e., $\exists t_{0}\in\mathcal{T}$, such that $\forall t\in\mathcal{T}$, it holds that $t_{0}\preceq t$, $t\preceq t_{a}$ and $t\preceq t_{b}$.
@@
Figure 2: A general model for coin flipping protocol using noiseless communication channels. The message spaces $X,Y$ and $X^{\prime},Y^{\prime}$ are of the same dimension, respectively.
In a WCF protocol, we say Alice wins if $c_{A}=c_{B}=0$ while Bob wins if $c_{A}=c_{B}=1$, which means Alice only wants to prevent Bob from biasing $c_{A}$ towards $1$ and Bob only wants to prevent Alice from biasing $c_{B}$ towards 00.
Notably, in the definition of weak coin flipping, players are not allowed to abort, whereas in strong coin flipping, aborting is required. This distinction arises because, in the weak coin flipping scenario, if a player detects malicious behavior, they can simply declare themselves the winner.
@@
Let $\epsilon\leq z$ and $z\in[0,\frac{1}{2}]$. A coin flipping protocol $\pi=(\pi_{A},\pi_{B})$ is a $z$-unbalanced $\epsilon$-biased WCF protocol (with stand-alone security), denoted by $\mathrm{WCF}(z,\epsilon)$, if the following conditions are satisfied:
@@
    For any converter $\alpha$ compatible with the inner interface with $\pi_{B}$, the composition $\alpha\pi_{B}$ outputs $c_{B}=0$ with a probability of
@@
    For any converter $\beta$ compatible with inner interface with $\pi_{A}$, the composition $\pi_{A}\beta$ outputs $c_{A}=1$ with a probability of
@@ 3 Weak coin flipping is not universally composably secure
This section shows that WCF cannot be universally composably secure in the AC framework. The result contains two parts. First, we consider a specific WCF resource derived from a simple but widely adapted SCF resource [30] and show that this specific WCF resource cannot be constructed from noiseless communication channel. Nevertheless, this result does not exclude the existence of other types of resources that can be constructed with, say, the protocol in [18]. Then our second result shows that, for any WCF protocol with enough stand-alone security, there does not exist any useful abstraction. Usefulness here means that the resource fully expresses the functionality of WCF and its interaction complexity is low.
@@ 3.1 Impossibility for a specific resource
Let $\epsilon\in[0,z]$ and $z\in[0,\frac{1}{2}]$. A $z$-unbalanced $\epsilon$-biased WCF recourse is characterized by a tuple $\mathcal{S}=(S,S_{A},S_{B})$, as is shown in Figure 3.
@@
    $S$ describes the case when both Alice and Bob are honest. In this case, both party obtains a random bit $c$ satisfying
@@
    $S_{A}$ describes the case when Alice is dishonest and Bob is honest. $S_{A}$ outputs a random bit $c^{\prime}$ subject to the distribution $(z,1-z)$ at $t_{0}^{\prime}$ and receives two bits $b^{\prime},p^{\prime}$ at $t_{1}^{\prime}$ at the left interface. Then it outputs a bit $c_{A}$ at the right interface. If $p^{\prime}=1$, then $c_{B}=1$, else the probability distribution of $c_{B}$ follows
    $\Pr[c_{B}=b^{\prime}]=\frac{\epsilon}{1-z},\quad\Pr[c_{B}=c^{\prime}]=1-\frac{\epsilon}{1-z}.$ (14)
    The causality condition is $t_{0}^{\prime}\prec t_{1}^{\prime}\prec t_{b}$.
@@
    $S_{B}$ works similarly as $S_{A}$. The causality condition is $t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}\prec t_{a}$.
$S$$(c,t_{a})$$(c,t_{b})$$S_{A}$$(c^{\prime},t_{0}^{\prime})$$(c_{B},t_{b})$$(b^{\prime},t_{1}^{\prime})$$S_{B}$$(c^{\prime\prime},t_{0}^{\prime\prime})$$(c_{A},t_{a})$$(b^{\prime\prime},t_{1}^{\prime\prime})$$(p^{\prime},t_{1}^{\prime})$$(p^{\prime\prime},t_{1}^{\prime\prime})$(a)(b)(c)
The resource in Definition 4 is a reasonable and simple WCF resource. The following theorem indicates that it cannot be constructed from noiseless communication channel.
@@
There does not exist a protocol $\pi=(\pi_{A},\pi_{B})$ which can construct an $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta<\frac{1}{3}z(1-z-\epsilon)$.
@@
Suppose that there is a protocol $\pi=(\pi_{A},\pi_{B})$ that constructs an ideal $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta$ (without loss of generality we assume $z\in[0,\frac{1}{2}]$). Combining the security conditions for Alice and Bob, respectively, followed by the triangle inequality, we obtain
@@
where $\sigma=\sigma_{B}\sigma_{A}$. Applying the correctness followed again by the triangle inequality, we proceed to
$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{3\delta}S.$ (16)
@@
Figure 4: The composition of two ideal WCF resources.
We now show that, however, there does not exist $\sigma$ such that $\delta<\frac{1}{12}$. Note that the best choice of causal order for $\sigma$ is $t_{0}^{\prime}\prec t_{1}^{\prime},t_{0}^{\prime}\prec t_{1}^{\prime\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}$ since it can maximize the use of information of $c^{\prime},c^{\prime\prime}$.
The most general operation of $\sigma$ is to apply a probabilistic map $\mathcal{M}_{\sigma}:(c^{\prime},c^{\prime\prime})\mapsto(b^{\prime},b^{\prime\prime},p^{\prime},p^{\prime\prime})$. Because $S$ must output $c_{A}=c_{B}=c$ while $S_{B}\sigma S_{A}$ might output $c_{A}\neq c_{B}$, a distinguisher $\mathcal{D}_{0}$ which only checks if $c_{A}=c_{B}$ to distinguish $S$ and $S_{B}\sigma S_{A}$ will provide a lower bound to the distinguishing advantage. That is, $\mathcal{D}_{0}(\cdot)$ outputs $1$ if $c_{A}=c_{B}$ and else 00. Consider the following map on three cases of the input.
@@
    $\Pr[c^{\prime}=c^{\prime\prime}=0]=z^{2}$ or $\Pr[c^{\prime}=c^{\prime\prime}=1]=(1-z)^{2}$. Because the bits are already the same, the box $\sigma$ outputs $p^{\prime}=0,b^{\prime}=c^{\prime},p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$ to keep $c_{A}$ and $c_{B}$ the same. We have $c_{A}=c_{B}=0$ or $c_{A}=c_{B}=1$.
@@
    $\Pr[c^{\prime}=0,c^{\prime\prime}=1]=z(1-z)$. The box $\sigma$ outputs $p^{\prime}=1,p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=1$. We have $c_{A}=c_{B}=1$.
@@
    $\Pr[c^{\prime}=1,c^{\prime\prime}=0]=z(1-z)$. The box $\sigma$ outputs $b^{\prime}=p^{\prime}=b^{\prime\prime}=p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=0$ with probability $\frac{\epsilon}{1-z}$.
This strategy of $\sigma$ is obviously the optimal one to maximize distinguishing probability under distinguisher $\mathcal{D}_{0}$, which is
$\Pr[c_{A}\neq c_{B}]=\Pr[c_{A}=1,c_{B}=0]+\Pr[c_{A}=0,c_{B}=1]=z(1-z-\epsilon).$ (17)
For any box $\sigma$, we have the lower bound on distinguishing advantage:
$d(S_{B}\sigma S_{A},S)\geq\left|\Pr[\mathcal{D}_{0}(S)=1]-\Pr[\mathcal{D}_{0}(S_{B}\sigma S_{A})=1]\right|\geq z(1-z-\epsilon).$ (18)
Together with the relation in Equation (16), we have
$\delta\geq\frac{1}{3}z(1-z-\epsilon).$ (19)
which shows that $\pi$ cannot construct $\mathcal{S}$ with a vanishing distinguishing advantage. ∎
By setting $\epsilon=0,z=\frac{1}{2}$, we conclude that there does not exist a protocol $\pi$ that constructs an ideal balanced unbiased WCF resource $\mathcal{S}$ with a distinguishing advantage $\delta<\frac{1}{12}$.

### v3

@@ 2.4 Weak coin flipping
Here, we formulate WCF protocols with causal boxes.
@@
A weak coin flipping protocol using noiseless communication channels is a pair of converters $\pi=(\pi_{A},\pi_{B})$ with compatible inner interfaces, where $\pi_{A}$ ($\pi_{B}$) has one output wire on the outer interface and only outputs a single classical bit at $t_{a}$ ($t_{b}$). The poset $\mathcal{T}$ of the protocol is bounded both from above and below, i.e., $\exists t_{0},t_{a},t_{b}\in\mathcal{T}$, such that $\forall t\in\mathcal{T}$, it holds that $t_{0}\preceq t$, $t\preceq t_{a}$ and $t\preceq t_{b}$.
@@
Figure 2: A general model for a coin flipping protocol using noiseless communication channels. The message spaces $X,Y$ and $X^{\prime},Y^{\prime}$ are of the same dimension, respectively.
In a WCF protocol, we say Alice wins if $c_{B}=0$, while Bob wins if $c_{A}=1$. In this case, an honest Bob only needs to prevent a cheating Alice from biasing $c_{B}$ towards 00 and an honest Alice only needs to prevent a cheating Bob from biasing $c_{A}$ towards $1$.
@@
Let $z\in[0,\frac{1}{2}]$ and $\epsilon\leq z$. A coin flipping protocol $\pi=(\pi_{A},\pi_{B})$ is a $z$-unbalanced $\epsilon$-biased WCF protocol (with stand-alone security), denoted by $\mathrm{WCF}(z,\epsilon)$, if the following conditions are satisfied:
@@
    For any converter $\alpha$ compatible with the inner interface of $\pi_{B}$, the composition $\alpha\pi_{B}$ outputs $c_{B}=0$ with a probability of
@@
    For any converter $\beta$ compatible with inner interface of $\pi_{A}$, the composition $\pi_{A}\beta$ outputs $c_{A}=1$ with a probability of
@@ 3 Weak coin flipping is not universally composably secure
This section shows that WCF cannot be universally composably secure in the AC framework. The result contains two parts. First, we consider a specific WCF resource derived from a simple but widely adapted SCF resource [30] and show that this specific WCF resource cannot be constructed from a noiseless communication channel. Nevertheless, this result does not exclude the existence of other more complicated resources that can be constructed with, say, the protocol in [18]. Then our second result shows that, for any WCF protocol with enough stand-alone security, one cannot abstract it into a useful black-box resource, where usefulness means that the resource captures the full functionality of WCF while maintains a low interaction complexity.
@@ 3.1 Impossibility for a specific resource
We consider a reasonable and simple WCF resource in Figure 3 which is largely motivated from the strong coin flipping resource described in [30, Section III]111 Although the resource had origins in Blum’s protocol [42] that constructs an unfair coin flip (where a party is allowed to abort on observing the outcome), it serves as an abstraction for a large family of SCF protocols. When using a similar protocol here, it is equivalent to a biased coin flip if the honest party declares themselves the winner (locally) whenever the dishonest party aborts [30]. Therefore, it is unnecessary to consider the case where players abort.. In the WCF task, a dishonest party wants to force only one of the two outcomes. To reflect this idea, we model the possible malicious operation as follows: (1) a cheating Alice (Bob) intercepts the random bit $c^{\prime}$ ($c^{\prime\prime}$) which ought to be the honest output in advance and (2) based on $c^{\prime}$ ($c^{\prime\prime}$), Alice (Bob) sends $b^{\prime}$ and $p^{\prime}$ ($b^{\prime\prime}$ and $p^{\prime\prime}$) to the resource to either resign by setting $p^{\prime}=1$ ($p^{\prime\prime}=1$) or replace $c^{\prime}$ with $b$ with probability $\frac{\epsilon}{1-z}$ ($\frac{\epsilon}{z}$) by setting $p^{\prime}=0$ ($p^{\prime\prime}=0$). This establishes Definition 4 as a tuple of ideal WCF resources that allows for a maximum bias of $\epsilon$ for different scenarios.
$S$$(c,t_{a})$$(c,t_{b})$$S_{A}$$(c^{\prime},t_{0}^{\prime})$$(c_{B},t_{b})$$((b^{\prime},p^{\prime}),t_{1}^{\prime})$$S_{B}$$(c^{\prime\prime},t_{0}^{\prime\prime})$$(c_{A},t_{a})$$((b^{\prime\prime},p^{\prime\prime}),t_{2}^{\prime\prime})$(a)(b)(c)
@@
Let $z\in[0,\frac{1}{2}]$ and $\epsilon\in[0,z]$. A $z$-unbalanced $\epsilon$-biased WCF resource is characterized by a tuple $\mathcal{S}=(S,S_{A},S_{B})$, as is shown in Figure 3.
@@
    $S$ describes the case when both Alice and Bob are honest. In this case, both parties obtains a random bit $c$ according to the probabilities
@@
    $S_{A}$ describes the case when Alice is dishonest and Bob is honest. $S_{A}$ outputs a random bit $c^{\prime}$ subject to the probability distribution $(z,1-z)$ at $t_{0}^{\prime}$ and receives two bits $b^{\prime},p^{\prime}$ at $t_{1}^{\prime}$ at the left interface. Then it outputs a bit $c_{B}$ at $t_{b}$ at the right interface. If $p^{\prime}=1$, then $c_{B}=1$, else $c_{B}=b^{\prime}$ with probability $\frac{\epsilon}{1-z}$ and $c_{B}=c^{\prime}$ with probability $1-\frac{\epsilon}{1-z}$. The causality condition is $t_{0}^{\prime}\prec t_{1}^{\prime}\prec t_{b}$.
@@
    $S_{B}$ describes the case when Alice is honest and Bob is dishonest. $S_{B}$ outputs a random bit $c^{\prime\prime}$ subject to the probability distribution $(z,1-z)$ at $t_{0}^{\prime\prime}$ and receives two bits $b^{\prime\prime},p^{\prime\prime}$ controlled by Bob at $t_{1}^{\prime\prime}$ at the right interface. Then it outputs a bit $c_{A}$ at $t_{a}$ at the left interface. If $p^{\prime\prime}=1$, then $c_{A}=0$, else $c_{A}=b^{\prime\prime}$ with probability $\frac{\epsilon}{z}$ and $c_{A}=c^{\prime\prime}$ with probability $1-\frac{\epsilon}{z}$. The causality condition is $t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}\prec t_{a}$.
To elaborate on the functionality when one party is malicious, consider an example where cheating Alice tries to bias the output $c_{B}$ of $S_{A}$ towards 00. For this purpose, she needs to set $p^{\prime}=0$ regardless of the value of $c^{\prime}$. Then the probability of her successfully forcing $c_{B}=0$ is
$\Pr[c_{B}=0]=\Pr[c^{\prime}=0]\cdot\Pr[c_{B}=0|c^{\prime}=0]+\Pr[c^{\prime}=1]\cdot\Pr[c_{B}=0|c^{\prime}=1].$
As $p^{\prime}=0$, the optimal strategy for Alice is to set $b^{\prime}=0$ since $c_{B}$ is either $c^{\prime}$ or $b^{\prime}$. If $c^{\prime}=0$, which happens with probability $\Pr[c^{\prime}=0]=z$, then $c_{B}=0$ for certain, i.e., $\Pr[c_{B}=0|c^{\prime}=0]=1$. If $c^{\prime}=1$, which happens with probability $\Pr[c^{\prime}=1]=1-z$, Alice’s conditional winning probability equals the probability that $c_{B}$ takes the value of $b^{\prime}$, i.e., $\Pr[c_{B}=0|c^{\prime}=1]=\frac{\epsilon}{1-z}$. We thus have
$\Pr[c_{B}=0]=\Pr[c^{\prime}=0]\cdot 1+\Pr[c^{\prime}=1]\cdot\frac{\epsilon}{1-z}=z+\epsilon.$ (14)
Similar arguments also hold for cheating Bob.
The following theorem indicates that the resource in Definition 4 cannot be constructed from noiseless communication channel.
@@
For $z\in[0,\frac{1}{2}]$ and $\epsilon\in[0,z]$, there does not exist a protocol $\pi=(\pi_{A},\pi_{B})$ which can construct an $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta<\frac{1}{3}\left(z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\}\right)$.
@@
Suppose that there exists a protocol $\pi=(\pi_{A},\pi_{B})$ that constructs an ideal $z$-unbalanced $\epsilon$-biased WCF resource $\mathcal{S}=(S,S_{A},S_{B})$ with a distinguishing advantage $\delta$ (without loss of generality, we assume $z\in[0,\frac{1}{2}]$). Combining the security conditions for Alice (6) and Bob (7) using triangle inequality, we obtain
@@
where $\sigma=\sigma_{B}\sigma_{A}$. Once again, combining the previous equation with the correctness condition (5) using triangle inequality, we get
$\exists\sigma\in\Sigma,S_{B}\sigma S_{A}\approx_{3\delta}S,$ (16)
or equivalently,
$\exists\sigma\in\Sigma,d(S_{B}\sigma S_{A},S)\leq 3\delta.$ (17)
@@
Figure 4: Illustration for Inequality (17).
We next establish a constant lower bound on such $\delta$ (constant in terms of $z$) by deriving a lower bound on the distinguishing advantage $d(S_{B}\sigma S_{A},S)$ for any $\sigma$, i.e., $\min_{\sigma}d(S_{B}\sigma S_{A},S)$. As the two systems $S_{B}\sigma S_{A}$ and $S$ in Figure 4 do not need any external input, $d(S_{B}\sigma S_{A},S)$ is simply the total variation distance between their output distributions, denoted by $P_{c_{A}c_{B}}$ and $P_{cc}$ respectively, i.e.,
$d(S_{B}\sigma S_{A},S)=\frac{1}{2}\|P_{c_{A}c_{B}}-P_{cc}\|_{1},$ (18)
with $P_{cc}(0,0)=z,P_{cc}(1,1)=1-z$.
The best choice of causal order for $\sigma$ is $t_{0}^{\prime}\prec t_{1}^{\prime},t_{0}^{\prime}\prec t_{1}^{\prime\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime},t_{0}^{\prime\prime}\prec t_{1}^{\prime\prime}$ since it can maximize the use of information of $c^{\prime},c^{\prime\prime}$. Then, the minimum distinguishing advantage can be bounded as follows:
$\min_{\sigma}d(S_{B}\sigma S_{A},S)$ $=\min_{\sigma}\frac{1}{2}\left(\Pr[c_{A}=0,c_{B}=1]+\Pr[c_{A}=1,c_{B}=0]\right.$ (19)
$\quad\quad+\left.\big{\lvert}\Pr[c_{A}=0,c_{B}=0]-z\big{\rvert}+\big{\lvert}\Pr[c_{A}=1,c_{B}=1]-1+z\big{\rvert}\right)$
$\geq\min_{\sigma}\left(\Pr[c_{A}=0,c_{B}=1]+\Pr[c_{A}=1,c_{B}=0]\right)$
$=\min_{\sigma}\Pr[c_{A}\neq c_{B}]$
$\geq\min_{\sigma}\Pr[c^{\prime}=1,c^{\prime\prime}=0]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]\enspace$
$\quad+\min_{\sigma}\Pr[c^{\prime}=0,c^{\prime\prime}=1]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=0,c^{\prime\prime}=1]\enspace$
$\quad+\min_{\sigma}\Pr[c^{\prime}=c^{\prime\prime}]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=c^{\prime\prime}].$
In the previous equation, the optimal distribution on $(b^{\prime},p^{\prime},b^{\prime\prime},p^{\prime\prime})$ determines the optimal $\sigma$ for each received value $c^{\prime},c^{\prime\prime}$. The optimal distributions for different choices of $c^{\prime}$ and $c^{\prime\prime}$ are discussed next.
@@
    If $c^{\prime}=c^{\prime\prime}=0$ or $c^{\prime}=c^{\prime\prime}=1$, then the box $\sigma$ outputs $p^{\prime}=0,b^{\prime}=c^{\prime},p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$ to keep $c_{A}=c^{\prime\prime}=c^{\prime}=c_{B}$, implying that $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=c^{\prime\prime}]=0$.
@@
    If $c^{\prime}=0,c^{\prime\prime}=1$, then the box $\sigma$ outputs $p^{\prime}=1,p^{\prime\prime}=0$ to force the outcome $c_{A}=c_{B}=1$, implying that $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=0,c^{\prime\prime}=1]=0$.
@@
    If $c^{\prime}=1,c^{\prime\prime}=0$, then we can assume the box $\sigma$ outputs $p^{\prime}=p^{\prime\prime}=0$ because setting $p^{\prime}=1$ ($p^{\prime\prime}=1$) has the same effect as setting $p^{\prime}=0,b^{\prime}=c^{\prime}$ ($p^{\prime\prime}=0,b^{\prime\prime}=c^{\prime\prime}$). Consider the event $c_{A}\neq c_{B}$ conditioned on different cases of $(b^{\prime},b^{\prime\prime})$:
    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=0,b^{\prime\prime}=0]=1-\frac{\epsilon}{1-z};$ (20)
    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=1,b^{\prime\prime}=0]=1;$
    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=0,b^{\prime\prime}=1]=1-\frac{\epsilon(1-2\epsilon)}{z(1-z)};$
    $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=1,b^{\prime\prime}=1]=1-\frac{\epsilon}{z}.$
    All possible cheating strategies are in the convex hull of the above four cases, then we have
    $\min_{\sigma}\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]$ (21)
    $=$ $\min_{P_{b^{\prime},b^{\prime\prime}}}\sum_{i^{\prime},i^{\prime\prime}\in\{0,1\}}P_{b^{\prime},b^{\prime\prime}}(i^{\prime},i^{\prime\prime})\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ (22)
    $=$ $\min_{i^{\prime},i^{\prime\prime}\in\{0,1\}}\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ (23)
    $=$ $\begin{cases}1-\frac{\epsilon(1-2\epsilon)}{z(1-z)},&\text{if }0\leq\epsilon<z/2\\ 1-\frac{\epsilon}{z},&\text{if }z/2<\epsilon\leq z\end{cases}.$ (24)
    where the conditional probabilities $\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0,b^{\prime}=i^{\prime},b^{\prime\prime}=i^{\prime\prime}]$ are given in Equation 20, and Equation 24 holds because $20\leq\epsilon\leq z\leq 1/$ 2.
By the above discussion, Equation 19 continues as
$\min_{\sigma}d(S_{B}\sigma S_{A},S)$ $\geq\min_{\sigma}\Pr[c^{\prime}=1,c^{\prime\prime}=0]\cdot\Pr[c_{A}\neq c_{B}|c^{\prime}=1,c^{\prime\prime}=0]$ (25)
$=z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\},$ (26)
where Equation 25 follows from Equation 19 and the implications of Item 1 and Item 2. Equation 26 simply follows from Equation 24. Finally, combining Equation 17 and Equation 26, we have
$\delta\geq\frac{1}{3}\left(z(1-z)-\epsilon+\epsilon\min\{2\epsilon,z\}\right),$ (27)
which implies that $\pi$ cannot construct $\mathcal{S}$ with a vanishing distinguishing advantage. ∎
Setting $\epsilon=0,z=\frac{1}{2}$, we conclude that there does not exist a protocol $\pi$ that constructs an ideal balanced unbiased WCF resource $\mathcal{S}$ with a distinguishing advantage $\delta<\frac{1}{12}$.